OpenStudy (anonymous):

(1/121-4x^6) help please

OpenStudy (unklerhaukus):

what is the question

OpenStudy (anonymous):

FACTOR COMPLETELY

OpenStudy (unklerhaukus):

\[(1/121-4x^6)\]\[=(1/11)^2-(2x^3)^2\]\[=(1/11-2x^3)(1/11+2x^3)\]

OpenStudy (anonymous):

THANkS so much!!